OpenStudy (anonymous):

The basic formula to calculate a student's GPA is the:

OpenStudy (anonymous):

are you asking how to do it??

OpenStudy (anonymous):

A=4 B=3 C=2 D=1 F=0 then add up all your classes then divide the number by how many classes you have example A,B,C,A,D,A,B 4+3+2+4+1+4+3=21 =21/ 7classes = 3.0 GPA